OpenStudy (anonymous):

use the sum/difference identities to evaluate exactly sin pi/12

OpenStudy (anonymous):

\[\pi/12*180/\pi\]=15 sin (45-30 )

OpenStudy (anonymous):

Use the difference formula: sin(x - y) = sin x cos y - sin y cos x x = pi/3 and y = pi/4; so x - y = pi/3 - pi/4 = pi/12

OpenStudy (phi):

first, write down the rule for sin(A-B) see http://www.intmath.com/analytic-trigonometry/2-sum-difference-angles.php

OpenStudy (anonymous):

sin45cos30-cos45sin30

OpenStudy (phi):

now replace with numbers

OpenStudy (anonymous):

\[\sqrt{4-4}/4\]

OpenStudy (anonymous):

right

OpenStudy (phi):

sin(15) won't be 0 did you use these values: sin 45 = sqr(2)/2 cos 30= sqr(3)/2 cos 45= sqr(2)/2 sin 30= ½

OpenStudy (anonymous):

\[\sqrt{2}-\sqrt{2}/4\]

OpenStudy (anonymous):

now it is right

OpenStudy (phi):

you should write down the expression first. You are going wrong somewheres, but I don't know where.

OpenStudy (anonymous):

\[\sqrt{6}-\sqrt{2}/4\]

OpenStudy (anonymous):

right

OpenStudy (phi):

sin45cos30-cos45sin30 sin 45 = sqr(2)/2 cos 30= sqr(3)/2 so the first term is sqr(2)/2 * sqr(3)/2 the 2nd term uses cos 45= sqr(2)/2 sin 30= ½ it is sqr(2)/2 * ½ and you now have sqr(2)/2 * sqr(3)/2 - sqr(2)/2 * ½ written more nicely: \[ \frac{\sqrt{2}}{2}\cdot \frac{\sqrt{3}}{2} - \frac{\sqrt{2}}{2} \cdot \frac{1}{2} \] multiply top times top and bottom times bottom you will have 4 as a common denominator, so you can put the top all over 4

OpenStudy (anonymous):

\[\sqrt{6}/4-\sqrt{2}/4\]

OpenStudy (anonymous):

right

OpenStudy (anonymous):

or \[\sqrt{6}-\sqrt{2}/4\]

OpenStudy (phi):

ok, but only if you put parens around the top \[ (\sqrt{6}-\sqrt{2})/4 = \frac{\sqrt{6}-\sqrt{2}}{4}\] if you leave out the parens, it means \[ \sqrt{6} - \frac{\sqrt{2}}{4} \] but yes, you have the answer

OpenStudy (anonymous):

thank you :)